Calculate the pH of the buffer that results from mixing 51.9 mL of a 0.298 M solution of HCHO2 and 16.2 mL of a 0.509 M solution of NaCHO2. The Ka value for HCHO2 is 1.8×10−4.

Express your answer to two decimal places.
Answer #1

millimoles of HCHO2 = 51.9 x 0.298 = 15.466

millimoles of NaCHO2 = 16.2 x 0.509 = 8.2458

pKa = 3.74

pH = pKa + log [salt / acid]

     = 3.74 + log [NaCHO2 / HCHO2]

pH = 3.74 + log [8.25 / 15.466]

pH = 3.47